OpenStudy (anonymous):

Use an angle sum identity to verify the identity cos 2 θ = 2 cos ^2 θ - 1

OpenStudy (anonymous):

cos(x+y)=cosxcosy-sinxsiny

OpenStudy (anonymous):

give x=y=theta...we get cos2theta=cos^2 theta-sin^2 theta

OpenStudy (anonymous):

\[\sin ^{2}\theta+\cos ^{2}\theta=1\]

OpenStudy (anonymous):

so sin^2 theta=1-cos^2 theta

OpenStudy (anonymous):

put it in our cos2theta eqn..we get cos^2theta-(1-cos^2 theta)

OpenStudy (anonymous):

this gives the answer 2cos^2 theta-1

OpenStudy (anonymous):

got it?@Aurelia0912

OpenStudy (anonymous):

yes thank you :)